OpenStudy (anonymous):

-28 ≤ -5t + -3 ≤ -18

OpenStudy (texaschic101):

-28 <= -5t - 3 <= -18 --- add 3 3 - 28 <= -5t - 3 + 3 <= -18 + 3 -- combine like terms -25 <= -5t <= - 15 --- divide by -5, and change the inequality signs -25/-5 >= -5t/-5 >= -15/-5 5 >= t >= 3 any questions ?

OpenStudy (anonymous):

|-2x + 3| = 15

OpenStudy (texaschic101):

| -2x + 3| = 15 (-2x + 3) = 15 -2x = 15 - 3 -2x = 12 x = -6 -(-2x + 3) = 15 2x - 3 = 15 2x = 15 + 3 2x = 18 x = 9 solution : x = -6 and x = 9